OpenStudy (anonymous):

Perform the indicated operation. Note that the denominators are the same. Simplify the result, if possible. \frac{y^{2}+3y-4}{y^{2}+4y+3} - \frac{y-1}{y^{2}+4y+3} So far i have got it down to (y-1)^2 (y+4)/(Y+3)(y+1)^2

OpenStudy (bibby):

\[\frac{y^{2}+3y-4}{y^{2}+4y+3} - \frac{y-1}{y^{2}+4y+3}\]

OpenStudy (anonymous):

So how do i simplify it now?

OpenStudy (bibby):

it might help if you realize that \[y^2+3y-4 = (y+4)(y-1)\]and that\[y^2+4y+3=(y+3)(y+1)\]

OpenStudy (anonymous):

Yep i can get to that step it just wont accept my answer on the homework site i dont know why

OpenStudy (bibby):

Notice the like denominators you can combine the terms so that you get \[\frac{ (y+4)(y-1)-(y-1) }{ (y+3)(y+1) }\]

OpenStudy (bibby):

TBH I'm not sure how to proceed